# Post-reduction inference for confidence sets of models

### v1

@@ 3.2. Distribution of the ancillary component
Notice that $R^{2}$ is the residual sum of squares $\hat{\varepsilon}^{{\rm T}}\hat{\varepsilon}$, which is part of the minimal sufficient statistic $S$ and has a distribution independent of $\theta$. It is ancillary for $\theta$ according to the more general of the definitions on page 38 of Barndorff-Nielsen and Cox (1994) and is therefore also available for the assessment of model adequacy via a version of (2.4) allowing for nuisance parameters.
Suppose initially that $\sigma^{2}$ and the mean subspace $\mathcal{X}$ are known, then $R^{2}$ carries all the information relevant for distinguishing, say, normal errors from Student-$t$ errors. This follows from the properties of elliptically symmetric distributions for which the generic form of the density function is
@@
Let $U$ be as defined in §3.1. Then $W=U^{{\rm T}}Y$ is a member of the elliptically symmetric class of distributions in $n-d_{\theta}$ dimensions, with zero mean and covariance $\Sigma_{W}=\sigma^{2}I_{n-d_{\theta}}=U^{{\rm T}}\Sigma U$ (e.g. Muirhead, 1982), and therefore has a stochastic representation (Cambanis et al., 1981, Theorem 1) $W\stackrel{{\scriptstyle d}}{{=}}Z\Sigma_{W}^{1/2}Q^{\prime}$, where =$\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution and $Q^{\prime}$ is uniformly distributed on the surface of the unit hypersphere in $\mathbb{R}^{n-d_{\theta}}$. The distribution of $Z$ with density function $f_{Z}$ determines the member of the elliptic class and is related to the generator $g$ for the standardised variable $\Sigma^{-1/2}_{W}W$ as (Fang et al., 1990)
@@
The random variable $Z^{2}$ thus carries information on the distribution of the errors when the mean model is known. Conversely, $Z^{2}=R^{2}/\sigma^{2}$ contains valuable information for assessment of the mean model under the assumption that the errors are Gaussian. The practical implementation raises further aspects, to which we return in §4.5.
@@ 4. An operational blueprint for Gaussian linear regression
4.1. Some initial observations
@@
From the data, we obtain one realisation $q^{o}$ of the random variable $Q$ from §3.1, which is uniformly distributed on the surface of the unit hypersphere $\mathcal{Q}(s^{o})$ in $\mathcal{X}^{\perp}$. A test based on $q^{o}$ alone has no power in an assessment of model adequacy, as all unit vectors on $\mathcal{Q}(s^{o})$ are compatible with uniformity of distribution. Although power in certain directions could be restored by consideration of alternative models for comparison, this leads to a test statistic essentially equivalent to the $F$-test, i.e. the likelihood ratio test for Gaussian models, and therefore back to the initial difficulties discussed in §1. Note, however, that even with two replicate observations of the random variable $Q$, there would be a degree of power against uniformity of distribution on $\mathcal{Q}(s^{o})$, this tending to increase with the number of replicate observations.
@@
This observation points to an initial suggestion for inference based on co-sufficient information partitioning. Information partitioning, while bearing some superficial similarity to sample splitting, differs in that no observational units are discarded in any of the inferential steps. Since all the data are used for both aspects of inference, reduction and model assessment, no information is relinquished in the reduction to $\hat{\mathcal{E}}$, and virtually no co-sufficient information is lost in the assessment of model adequacy. As it turns out, the simplest partitioning approach, while conceptually revealing, is ineffectual for reasons given in §4.2. We propose in §4.3 a more elaborate reformulation, mirroring to some extent the approach of Rasines and Young (2023) in a different context.
4.2. Preliminary insight: power and naïve data-partitioning
From an encompassing model $\hat{\mathcal{E}}$ computed using all available data, we consider all submodels $\mathcal{E}_{m}$ of reasonable size, seeking to assess each one for compatibility with the data. From $k$ disjoint subsamples of size $n_{k}=n/k$, we consider initially $Q^{(1)},\ldots,Q^{(k)}$, constructed as above but based on disjoint subsamples $(Y^{(1)},X^{(1)}),\ldots,(Y^{(k)},X^{(k)})$.
$\frac{\exp\{-(a+w^{{\rm T}}w)/2\sigma^{2}\}}{(2\pi\sigma^{2})^{(n-d_{\theta})/2}}\exp\Bigl{(}-\frac{w^{{\rm T}}U^{{\rm T}}Z\lambda}{\sigma^{2}}\Bigr{)}dw,$ (4.1)
In the absence of a closed form for the marginal distribution of $Q$, Appendix §A.1 establishes the approximation
$f_{Q}(q)d\text{Vol}^{n-d}(q)\simeq\frac{\Gamma\bigl{(}\frac{n-d_{\theta}}{2}\bigr{)}}{2\pi^{(n-d_{\theta})/2}}\exp\Bigl{(}-\frac{a}{2\sigma^{2}}\Bigr{)}\exp\biggl{\{}\frac{(n-d_{\theta})q^{{\rm T}}U^{{\rm T}}Z\lambda}{\sigma\sqrt{2}}\biggr{\}}d\text{Vol}^{n-d}(q),$ (4.2)
which we write as
Approximation (4.2), although obtained under the notional limiting operation $(n-d_{\theta})\rightarrow\infty$, is essentially exact, the only error coming from Stirling’s approximation.
The first multiplicative factor on the right hand side of (4.2) is the density function of $Q$ under the null hypothesis. Thus, power is low when $h(q)$ is close to 1 for all $q$. This occurs when $Z\lambda$ is close to zero or when $Z$ lies close to the column span of $X$. The conclusion agrees qualitatively with the power analysis for the likelihood-ratio test in the absence of any dependencies from reduction, as derived in §4.2 of Lewis and Battey (2025) using a limiting argument in Le Cam-type contiguous alternatives.
Consider now a test of uniformity on the unit hypersphere in $\mathbb{R}^{n_{k}-d_{\theta}}$ based on the partitioned data, discarding the orientation of each $\mathcal{X}^{(i)\perp}$ within $\mathbb{R}^{n_{k}}$; $k$ now plays the role of sample size for the null uniform samples. Equation (4.2) suggests that the non-null distribution of $Q^{(i)}$ is well approximated by a von Mises-Fisher distribution with mean direction $m^{(i)}=U^{(i)T}Z^{(i)}\lambda$ and dispersion parameter $\kappa^{(i)}=\sigma^{-1}2^{-1/2}(n_{k}-d_{\theta})\|U^{(i)T}Z^{(i)}\lambda\|$. Since the mean vectors $m^{(i)}$ depend on how the data are partitioned, and because these themselves are in general numerically indistinguishable from uniformly distributed random vectors on the unit hypersphere in $\mathbb{R}^{n_{k}-d_{\theta}}$ when standardised to unit length, the distribution of the test statistic when the assumed model is violated is hard to distinguish from that under the model postulated under the null hypothesis. Put differently, by ignoring the orientation of the embedding manifold $\mathcal{X}^{(i)\perp}$, the analysis is made too unconditional.
@@ 4.3. Co-sufficient information partitioning through randomisation
$[\tilde{Y}^{(1)}\cdots\tilde{Y}^{(k)}]=[Y\;L]\Gamma,$ (4.3)
@@
$\Gamma=\begin{bmatrix}1&1&1&\cdots&1\\ a_{1}&-b_{1}&-b_{1}&\cdots&-b_{1}\\ 0&a_{2}&-b_{2}&\cdots&-b_{2}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&a_{k-1}&-b_{k-1}\end{bmatrix}.$ (4.4)
@@
$\begin{split}\tilde{a}_{1}=&\;\sqrt{k-1};\hskip 91.04872pt\tilde{b}_{1}=\frac{1}{\tilde{a}_{1}};\\ \tilde{a}_{i+1}=&\;\Bigl{(}k-1-\sum_{j=1}^{i}\tilde{b}_{j}^{2}\Bigr{)}^{1/2};\hskip 27.88358pt\tilde{b}_{i+1}=\frac{1+\sum_{j=1}^{i}\tilde{b}_{j}^{2}}{\tilde{a}_{i+1}}.\end{split}$ (4.5)
@@
A proof is provided in Appendix A.2. The above construction can be generalised to allow for any known positive-definite covariance matrix $\Sigma$ in the distribution of $Y$ by adjusting $\Gamma$ appropriately.
@@
This decomposition of the original response vector yields independent replicates of $Q$, constructed as $\tilde{Q}^{(i)}=U^{{\rm T}}\tilde{Y}^{(i)}/\|U^{{\rm T}}\tilde{Y}^{(i)}\|$ for $i=1,\ldots,k$, where $U$ is that of §3.1. Under correct submodel specification, $\mu=X\theta$, these $k$ replicates are independent and uniformly distributed on the unit hypersphere in $\mathcal{X}^{\perp}$.
@@
For an actionable version of this approach, a reliable estimate of $\sigma^{2}$ is needed. The classical estimator, constructed from the residual sum of squares of the linear model including all covariates in the encompassing set $\hat{\mathcal{E}}$, is consistent provided that $n-|\hat{\mathcal{E}}|\to\infty$. Since $\hat{\mathcal{E}}$ tends to overfit to the data in small samples, it typically underestimates $\sigma^{2}$ unless the sample size is large. Small-sample bias is reduced through the use of the refitted cross-fitting procedure of Fan et al. (2012), who provided statistical guarantees, or the minor adaptation of this estimator proposed below, whose distribution theory is simple and convenient for the Gaussian linear model.
@@
The estimator described by Fan et al (2012) partitions the outcome and covariate data into two disjoint subsets, $(Y^{(1)},X^{(1)})$ and $(Y^{(2)},X^{(2)})$, each of size $n/2$. Fan et al. (2012) proposed to fit a conservative variable selection procedure within each subset, such that the true set $\mathcal{E}^{*}$ is contained in both selected sets ^$\hat{\mathcal{E}}^{(1)}$ and ^$\hat{\mathcal{E}}^{(2)}$ with high probability. Let $\mathcal{X}^{(1)}_{2}$ be the space spanned by the columns of $X^{(1)}$ indexed by ^$\hat{\mathcal{E}}^{(2)}$ and let $\mathcal{X}^{(2)}_{1}$ be the space spanned by the columns of $X^{(2)}$ indexed by ^$\hat{\mathcal{E}}^{(1)}$. The projection matrices onto these two subspaces are written $P^{(1)}_{2}$ and $P^{(2)}_{1}$. The refitted cross-validation estimator $\hat{\sigma}^{2}_{\text{rcv}}$ of $\sigma^{2}$ is then given by the average of the two independent estimators
$\hat{\sigma}_{1}^{2}=\frac{\|P^{(1)}_{2}Y^{(1)}\|_{2}^{2}}{n/2-|\hat{\mathcal{E}}^{(2)}|},\quad\quad\hat{\sigma}_{2}^{2}=\frac{{\|P^{(2)}_{1}Y^{(2)}\|_{2}^{2}}}{n/2-|\hat{\mathcal{E}}^{(1)}|}.$
@@
Empirically, the best approach computes this estimate, or the one below, using only $[\gamma n]$ observations for some $10.5<\gamma<$ 1, so that $n$ in the previous display is replaced by $[\gamma n]$. This further limits the dependence between $\mathop{\hat{\sigma}^{2}_{\mathrm{rcv}}}$ and $U^{T}Y$, resulting in a better calibration of the test statistic under the null distribution. In simulations, we found that $\gamma=0.6$ was a good choice across a broad range of scenarios.
@@
$\hat{\sigma}^{2}_{\text{mrcv}}:=\frac{df_{1}\hat{\sigma}^{2}_{1}+df_{2}\hat{\sigma}^{2}_{2}}{df_{1}+df_{2}}.$ (4.6)
@@
Let $\nu=df_{1}+df_{2}$. Conditional on |^$|\hat{\mathcal{E}}^{(1)}|$ and |^$|\hat{\mathcal{E}}^{(2)}|$, and on the containments ^$\mathcal{E}^{*}\subseteq\hat{\mathcal{E}}^{(1)}$ and ^$\mathcal{E}^{*}\subseteq\hat{\mathcal{E}}^{(2)}$,
$\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}\sim\sigma^{2}\frac{\chi^{2}_{\nu}}{\nu}$
and
$\mathop{\mathrm{pr}}\bigl{(}\bigl{|}\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}/\sigma^{2}-1\bigr{|}>\delta\bigr{)}\leq 2\exp\bigl{\{}-\textstyle{\frac{1}{2}}\bigl{[}(\nu+\delta+1)+\nu\log(\nu/(\delta+1)\bigr{]}\bigr{\}}.$
Let $\hat{Y}^{(1)},\ldots,\hat{Y}^{(k)}$ be the $k$ columns of the $n\times k$ matrix obtained by replacing $\Gamma$ in (4.3) by $\hat{\Gamma}$, the realisable version of $\Gamma$ with $\hat{\sigma}$ in place of $\sigma$. Our recommended choice of estimator $\hat{\sigma}$ is $\hat{\sigma}_{\text{mrcv}}$ computed on a portion $[0.6n]$ of the data. With $\hat{Q}^{(i)}=U^{{\rm T}}\hat{Y}^{(i)}/\|U^{{\rm T}}\hat{Y}^{(i)}\|$, Proposition 4.3 quantifies the maximum Euclidean distance between $\hat{Q}^{(1)},\ldots,\hat{Q}^{(k)}$ and the idealised versions $\tilde{Q}^{(1)},\ldots,\tilde{Q}^{(k)}$ that treat $\sigma$ as known. The conclusion is in terms of $k$ and the error rate $a_{n}$ of the estimator $\hat{\sigma}$ of $\sigma$.
It should be kept in mind that, although the notation does not reflect this, the construction of both $\hat{Q}^{(1)},\ldots,\hat{Q}^{(k)}$ and $\tilde{Q}^{(1)},\ldots,\tilde{Q}^{(k)}$ depends on a postulated model $\mathcal{E}_{m}$, inducing a mean subspace, $\mathcal{X}_{m}$ say.
@@
In contrast to the simple data-partitioning scheme considered in §4.2, the replicates of $Q$ obtained through randomisation via (4.3) are approximately independent and identically distributed under mean models other than the postulated one $\mu=X\theta$. As a consequence, this decomposition is more amenable than that of §4.2 to standard testing procedures. While there are many tests of uniformity on a sphere, we focus on the high-dimensional Rayleigh test, as its properties are guaranteed under a high-dimensional regime that matches our requirements, and efficient implementations are available in standard software.
@@
$R_{n}:=\frac{\sqrt{2(n-d_{\theta})}}{k}\sum_{1\leq i<j\leq k}\left<\frac{U^{{\rm T}}\tilde{Y}^{(i)}}{\lVert U^{{\rm T}}\tilde{Y}^{(i)}\rVert_{2}},\frac{U^{{\rm T}}\tilde{Y}^{(j)}}{\lVert U^{T}\tilde{Y}^{(j)}\rVert_{2}}\right>\rightarrow_{d}N(0,1),$ (4.7)
or rather, its analogue with $\hat{Y}^{(i)}$ in place of $\tilde{Y}^{(i)}$. This test rejects for large values of $R_{n}$, calibrated against the asymptotic null distribution. This is consistent with a direction-free alternative hypothesis: if $\mu\notin\mathcal{X}$, then $\tilde{Q}^{(1)},\ldots,\tilde{Q}^{(k)}$ are approximately a random sample from projected Gaussian with non-zero mean, so small angles among them are indicative of model misspecification. The distributional convergence was ascertained by Paindaveine and Verdebout (2016) in a more general setting than that needed here. Since, under the null hypothesis that the postulated mean model is correct, $U^{{\rm T}}\tilde{Y}^{(i)}$ is normally distributed of mean zero and variance $k\sigma^{2}I_{n-d_{\theta}}$, exact distribution theory for $R_{n}$ can in principle be derived from the calculations of Fisher (1915). We have
$\frac{U^{{\rm T}}\tilde{Y}^{(i)}}{\lVert U^{{\rm T}}\tilde{Y}^{(i)}\rVert_{2}}=\frac{\sigma^{-1}k^{-1/2}U^{{\rm T}}\tilde{Y}^{(i)}}{\lVert\sigma^{-1}k^{-1/2}U^{{\rm T}}\tilde{Y}^{(i)}\rVert_{2}},$ (4.8)
@@
$\frac{U^{{\rm T}}\tilde{Y}^{(i)}}{\lVert U^{{\rm T}}\tilde{Y}^{(i)}\rVert_{2}},\frac{U^{{\rm T}}\tilde{Y}^{(j)}}{\lVert U^{T}\tilde{Y}^{(j)}\rVert_{2}}\right>=\cos\left(\angle\left(\frac{U^{{\rm T}}\tilde{Y}^{(i)}}{k^{1/2}\sigma},\frac{U^{{\rm T}}\tilde{Y}^{(j)}}{k^{1/2}\sigma}\right)\right)=:\bar{R},$

### v2

@@ 4.2. Distribution of the ancillary component
Notice that $R^{2}$ is the residual sum of squares $\hat{\varepsilon}^{{\rm T}}\hat{\varepsilon}$, which is part of the minimal sufficient statistic $S$ and has a distribution independent of $\theta$. It is ancillary for $\theta$ according to the more general definitions on page 38 of Barndorff-Nielsen and Cox (1994) and is therefore also available for the assessment of model adequacy via a version of (3.4) allowing for nuisance parameters.
Suppose initially that $\sigma^{2}$ and the mean subspace $\mathcal{X}$ are known, then $R^{2}$ carries all the information relevant for distinguishing, say, normal errors from Student-$t$ errors. This follows from the properties of elliptically symmetric distributions, for which the generic form of the density function is
@@
Let $U$ be as defined in §4.1. Then $W=U^{{\rm T}}Y$ is a member of the elliptically symmetric class of distributions in $n-d_{\theta}$ dimensions, with zero mean and covariance $\Sigma_{W}=\sigma^{2}I_{n-d_{\theta}}=U^{{\rm T}}\Sigma U$ (e.g. Muirhead, 1982), and therefore has a stochastic representation (Cambanis et al., 1981, Theorem 1) $W\stackrel{{\scriptstyle d}}{{=}}Z\Sigma_{W}^{1/2}Q^{\prime}$, where $\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution and $Q^{\prime}$ is uniformly distributed on the surface of the unit hypersphere in $\mathbb{R}^{n-d_{\theta}}$. The distribution of $Z$ with density function $f_{Z}$ determines the member of the elliptic class and is related to the generator $g$ for the standardised variable $\Sigma^{-1/2}_{W}W$ as (Fang et al., 1990)
@@
The random variable $Z^{2}$ thus carries information on the distribution of the errors when the mean model is known. Conversely, $Z^{2}=R^{2}/\sigma^{2}$ contains valuable information for assessment of the mean model under the assumption that the errors are Gaussian.
@@ 5. An operational blueprint for Gaussian linear regression
5.1. Co-sufficiency and confidence sets of models: difficulties
Sections 3 and 4 covered, respectively, general principles and the corresponding calculations of the normal-theory linear model. These sections were detached from the problem of post-reducton inference for confidence sets of models introduced in §1, to which we now return.
From an encompassing model $\hat{\mathcal{E}}$ computed using all available data, we consider all submodels $\mathcal{E}_{m}$ of reasonable size, seeking to assess each one for compatibility with the data. Thus for each submodel $\mathcal{E}_{m}$, we can construct one realisation $q^{o}$ of the random variable $Q$ from §4.1. The dependence of $Q$ on the model $\mathcal{E}_{m}$ is not made explicit in the notation; nor is that of the corresponding null space $\mathcal{X}^{\perp}$ or the unit hypersphere $\mathcal{Q}(s^{o})$ on which the distribution of $Q$ is supported. Under the correct model $\mathcal{E}^{*}$, the outcome $Y$ projected onto the corresponding $\mathcal{Q}(s^{o})$ is uniformly distributed on its surface.
@@ 5.1. Co-sufficiency and confidence sets of models: difficulties
The difficulty is that a test based on $q^{o}$ alone has no power in an assessment of model adequacy, as any single unit vectors on $\mathcal{Q}(s^{o})$ is compatible with uniformity of distribution. Although power in certain directions could be restored by consideration of alternative models for comparison, this leads to a test statistic essentially equivalent to the $F$-test, i.e. the likelihood ratio test for Gaussian models, and therefore back to the initial difficulties discussed in §1. Note, however, that even with two replicate observations of the random variable $Q$, there would be a degree of power against uniformity of distribution on $\mathcal{Q}(s^{o})$, this tending to increase with the number of replicate observations.
@@
This observation points to an initial suggestion for inference based on co-sufficient information partitioning. Information partitioning, while bearing some superficial similarity to sample splitting, differs in that no observational units are discarded in any of the inferential steps. Since all the data are used for both aspects of inference, reduction and model assessment, no information is relinquished in the reduction to $\hat{\mathcal{E}}$, and virtually no co-sufficient information is lost in the assessment of model adequacy. As it turns out, the simplest partitioning approach, while conceptually revealing, is ineffectual for reasons given in Appendix A. Instead, we propose in §5.2 a more elaborate reformulation, adapting and extending the approach of Rasines and Young (2023), introduced in that work for a different purpose.
@@ 5.2. Co-sufficient information partitioning through randomisation
(5.1) $\tilde{Y}^{(1)}\cdots\tilde{Y}^{(k)}]=[Y\;L]\Gamma,$
@@
(5.2) $\Gamma=\begin{bmatrix}1&1&1&\cdots&1\\ a_{1}&-b_{1}&-b_{1}&\cdots&-b_{1}\\ 0&a_{2}&-b_{2}&\cdots&-b_{2}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&a_{k-1}&-b_{k-1}\end{bmatrix}$.
@@
(5.3) $\begin{split}\tilde{a}_{1}=&\;\sqrt{k-1};\hskip 91.04872pt\tilde{b}_{1}=\frac{1}{\tilde{a}_{1}};\\ \tilde{a}_{i+1}=&\;\Bigl(k-1-\sum_{j=1}^{i}\tilde{b}_{j}^{2}\Bigr)^{1/2};\hskip 27.88358pt\tilde{b}_{i+1}=\frac{1+\sum_{j=1}^{i}\tilde{b}_{j}^{2}}{\tilde{a}_{i+1}}.\end{split}$
@@
A proof is provided in Appendix B.2. The above construction can be generalised to allow for any known positive-definite covariance matrix $\Sigma$ in the distribution of $Y$ by adjusting $\Gamma$ appropriately.
The randomisation scheme is a direct generalisation of the construction introduced by Rasines and Young (2023) in a different context. In particular, Rasines and Young (2023) decompose a random vector $Y\sim N(\mu,\sigma^{2}I_{n})$ into two vectors of the same size
$Y^{(1)}=$ $Y+\gamma\sigma L$ $\sim N(\mu,(1+\gamma^{2})\sigma^{2}I_{n}),$
$Y^{(2)}=$ $\;\,Y-\sigma L/\gamma$ $\sim N(\mu,(1+\gamma^{-2})\sigma^{2}I_{n}),$
where $L\sim N(0,1)$ and $\gamma>0$ calibrates the variance allocation. On applying this argument recursively k1k-1 times, we have, with $\gamma_{1}=(k-1)^{1/2}$, a decomposition of $Y$ into $Y^{(1)}\sim N(\mu,k\sigma^{2}I_{n})$ and $V_{1}\sim N(\mu,(k/k-1)\sigma^{2}I_{n})$, and with $\gamma_{j}=(k-j)^{1/2}$ a decomposition of $V_{j-1}$ into $Y^{(j)}\sim N(\mu,k\sigma^{2}I_{n})$ and $V_{j}\sim N(\mu,(k/k-j)\sigma^{2}I_{n})$. The matrix $\Gamma$ emerges on writing $Y^{(i)}$ in terms of $Y$ and the Gaussian errors, and collecting coefficients. The same construction can be obtained by an application of Proposition 2 in Dharamshi et al. (2025) but the explicit form of $\Gamma$ that achieves a decomposition into equal-variance components is new.
@@
From the pseudo-replicates of the response vector, we obtain independent replicates of $Q$, constructed for any given model as $\tilde{Q}^{(i)}=U^{{\rm T}}\tilde{Y}^{(i)}/\|U^{{\rm T}}\tilde{Y}^{(i)}\|$ for $i=1,\ldots,k$, where $U$ is that of §4.1. Under correct submodel specification, $\mu=X\theta$, these $k$ replicates are independent and uniformly distributed on the unit hypersphere in $\mathcal{X}^{\perp}$.
If $\sigma^{2}$ is known or precisely and accurately estimated, then a larger choice of $k$ is intuitively better, as this plays the role of sample size for model assessment, determining the number of observations on the sphere. Since an actionable version of the randomisation approach requires estimation of $\sigma^{2}$, which induces some asymptotically negligible dependence between pseudo replicates, small $k$ is better at small sample sizes, a point to which we return below.
@@
The classical estimator, constructed from the residual sum of squares of the linear model including all covariates in the encompassing set $\hat{\mathcal{E}}$, is consistent provided that n|^|n-$\hat{\mathcal{E}}|\to\infty$. Since $\hat{\mathcal{E}}$ tends to overfit to the data in small samples, it typically underestimates $\sigma^{2}$ unless the sample size is large. Small-sample bias is reduced through the use of the refitted cross-fitting procedure of Fan et al. (2012), who provided statistical guarantees. Below we propose a minor adaptation of this estimator, whose distribution theory is simple modulo the necessary conditioning arguments, and convenient for the Gaussian linear model.
@@
The estimator described by Fan et al (2012) partitions the outcome and covariate data into two disjoint subsets, $(Y^{(1)},X^{(1)})$ and $(Y^{(2)},X^{(2)})$, each of size $n/2$. Fan et al. (2012) proposed to fit a conservative variable selection procedure within each subset, such that the true set $\mathcal{E}^{*}$ is contained in both selected sets $\hat{\mathcal{E}}^{(1)}$ and $\hat{\mathcal{E}}^{(2)}$ with high probability. Let $\mathcal{X}^{(1)}_{2}$ be the space spanned by the columns of $X^{(1)}$ indexed by $\hat{\mathcal{E}}^{(2)}$ and let $\mathcal{X}^{(2)}_{1}$ be the space spanned by the columns of $X^{(2)}$ indexed by $\hat{\mathcal{E}}^{(1)}$. The projection matrices onto these two subspaces are written $P^{(1)}_{2}$ and $P^{(2)}_{1}$. The refitted cross-validation estimator $\hat{\sigma}^{2}_{\text{rcv}}$ of $\sigma^{2}$ is then given by the average of the two independent estimators
$\hat{\sigma}_{1}^{2}=\frac{\|(I-P^{(1)}_{2})Y^{(1)}\|_{2}^{2}}{n/2-|\hat{\mathcal{E}}^{(2)}|},\quad\quad\hat{\sigma}_{2}^{2}=\frac{{\|(I-P^{(2)}_{1})Y^{(2)}\|_{2}^{2}}}{n/2-|\hat{\mathcal{E}}^{(1)}|}.$
@@
Empirically, the best approach computes this estimate, or the one below, using only $\gamma n]$ observations for some $0.5<\gamma<1$, so that $n$ in the previous display is replaced by $\gamma n]$. This further limits the dependence between $\mathop{\hat{\sigma}^{2}_{\mathrm{rcv}}}$ and $U^{T}Y$, resulting in a better calibration of the test statistic under the null distribution. In simulations, we found that $\gamma=0.6$ was a good choice across a broad range of scenarios.
@@
(5.4) $\hat{\sigma}^{2}_{\text{mrcv}}:=\frac{df_{1}\hat{\sigma}^{2}_{1}+df_{2}\hat{\sigma}^{2}_{2}}{df_{1}+df_{2}}.$
Proposition 5.2 establishes a distributional approximation to the distribution of this estimator when the sizes of the compassing models, $|\hat{\mathcal{E}}^{(1)}|$ and $|\hat{\mathcal{E}}^{(2)}|$, are fixed independently of the data. Although in the case of fixed $|\hat{\mathcal{E}}^{(1)}|$ and $|\hat{\mathcal{E}}^{(2)}|$ they would typically be chosen equal so that $\mathop{\hat{\sigma}^{2}_{\mathrm{mrcv}}}$ reduces to $\mathop{\hat{\sigma}^{2}_{\mathrm{rcv}}}$, simulations (not reported) show that the distributional approximation given in Proposition 5.2 is almost exact at moderate sample sizes even when $|\hat{\mathcal{E}}^{(1)}|$ and $|\hat{\mathcal{E}}^{(2)}|$ are data-dependent.
@@
Let $\nu=df_{1}+df_{2}$. Let $|\hat{\mathcal{E}}^{(1)}|$ and $|\hat{\mathcal{E}}^{(2)}|$ be fixed independently of the data and satisfy $|\hat{\mathcal{E}}^{(1)}|+|\hat{\mathcal{E}}^{(2)}|=o(n)$. Assume that for any set $E$ such that $E\subseteq\{1,\ldots,p\}$ and $|E|=|\hat{\mathcal{E}}^{(-j)}|$, $X^{(j)}_{E}$ has full column rank for $j=1,2$, where the notation j-j denotes the complementary index. Assume further that $\mathop{\mathrm{pr}}(\mathcal{E}^{*}\not\subseteq\hat{\mathcal{E}}^{(1)}\cap\hat{\mathcal{E}}^{(2)})=O(n^{-1/2})$. Then,
$\hat{\sigma}^{2}\stackrel{{\scriptstyle d}}{{=}}\sigma^{2}\,\frac{\chi^{2}_{\nu}}{\nu}+O_{p}(n^{-1/2}),\quad\text{as }n\to\infty,$
where $\chi^{2}_{\nu}$ is a chi-square random variable with $\nu$ degrees of freedom.
Let $\hat{Y}^{(1)},\ldots,\hat{Y}^{(k)}$ be the $k$ columns of the $n\times k$ matrix obtained by replacing $\Gamma$ in (5.1) by $\hat{\Gamma}$, the realisable version of $\Gamma$ with $\hat{\sigma}$ in place of $\sigma$. A recommendation is $\hat{\sigma}_{\text{mrcv}}$ computed on a portion $[0.6n]$ of the data. With $\hat{Q}^{(i)}=U^{{\rm T}}\hat{Y}^{(i)}/\|U^{{\rm T}}\hat{Y}^{(i)}\|$, Proposition 5.3 quantifies the maximum Euclidean distance between $\hat{Q}^{(1)},\ldots,\hat{Q}^{(k)}$ and the idealised versions $\tilde{Q}^{(1)},\ldots,\tilde{Q}^{(k)}$ that treat $\sigma$ as known. The conclusion is in terms of $k$ and the error rate $a_{n}$ of the estimator $\hat{\sigma}$ of $\sigma$. It should be kept in mind that, although the notation does not reflect this, the construction of both $\hat{Q}^{(1)},\ldots,\hat{Q}^{(k)}$ and $\tilde{Q}^{(1)},\ldots,\tilde{Q}^{(k)}$ depends on a postulated model $\mathcal{E}_{m}$, inducing a mean subspace, $\mathcal{X}_{m}$ say.
@@
In contrast to the simple data-partitioning scheme considered in § A, the replicates of $Q$ obtained through randomisation via (5.1) are approximately independent and identically distributed under mean models other than the postulated one $\mu=X\theta$. As a consequence, this decomposition is more amenable than that of § A to standard testing procedures. While there are many tests of uniformity on a sphere, we focus on the high-dimensional Rayleigh test, as its properties are guaranteed under a high-dimensional regime that matches our requirements, and efficient implementations are available in standard software.
@@
(5.5) $R_{n}:=\frac{\sqrt{2(n-d_{\theta})}}{k}\sum_{1\leq i<j\leq k}\left<\frac{U^{{\rm T}}\tilde{Y}^{(i)}}{\lVert U^{{\rm T}}\tilde{Y}^{(i)}\rVert_{2}},\frac{U^{{\rm T}}\tilde{Y}^{(j)}}{\lVert U^{T}\tilde{Y}^{(j)}\rVert_{2}}\right>\rightarrow_{d}N(0,1),$
or rather, its analogue with $\hat{Y}^{(i)}$ in place of $\tilde{Y}^{(i)}$. This test rejects for large values of $R_{n}$, calibrated against the asymptotic null distribution. This is consistent with a direction-free alternative hypothesis: if $\mu\notin\mathcal{X}$, then $\tilde{Q}^{(1)},\ldots,\tilde{Q}^{(k)}$ are approximately a random sample from projected Gaussian with non-zero mean, so small angles among them are indicative of model misspecification. The distributional convergence was ascertained by Paindaveine and Verdebout (2016) in a more general setting than that needed here. Since, under the null hypothesis that the postulated mean model is correct, $U^{{\rm T}}\tilde{Y}^{(i)}$ is normally distributed with mean zero and variance $k\sigma^{2}I_{n-d_{\theta}}$, exact distribution theory for $R_{n}$ can in principle be derived from the calculations of Fisher (1915). We have
(5.6) $\frac{U^{{\rm T}}\tilde{Y}^{(i)}}{\lVert U^{{\rm T}}\tilde{Y}^{(i)}\rVert_{2}}=\frac{\sigma^{-1}k^{-1/2}U^{{\rm T}}\tilde{Y}^{(i)}}{\lVert\sigma^{-1}k^{-1/2}U^{{\rm T}}\tilde{Y}^{(i)}\rVert_{2}},$
@@
$\left<\frac{U^{{\rm T}}\tilde{Y}^{(i)}}{\lVert U^{{\rm T}}\tilde{Y}^{(i)}\rVert_{2}},\frac{U^{{\rm T}}\tilde{Y}^{(j)}}{\lVert U^{T}\tilde{Y}^{(j)}\rVert_{2}}\right>=\cos\left(\angle\left(\frac{U^{{\rm T}}\tilde{Y}^{(i)}}{k^{1/2}\sigma},\frac{U^{{\rm T}}\tilde{Y}^{(j)}}{k^{1/2}\sigma}\right)\right)=:\bar{R},$